I just want to check if my answer is correct.
Problem:the center of the circle is at (-3,-2). If a chord of length 4 is bisected at (3,1), find the length of the radius.
My answer is: √21

Answer by rothauserc(4718)

let the distance between the two given points be d, r is radius, then we have
d^2 + 2^2 = r^2
d = square root( (-3 -3)^2 + (-2 -1)^2 )
d = square root ( 36 + 9) and
d^2 = 45, therefore
r^2 = 45 + 4 = 49
r = 7